# Quasi-randomization tests for network interference

### v1

@@ 2.3   Null hypothesis of no spillover and conditional randomization test
$\forall i\in[N]\quad Y_{i}(Z,G)=Y_{i}(Z^{\prime},G^{\prime})\qquad$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
@@
$\forall i\in[N]\quad Y^{2}(Z_{\mathcal{N}_{G}(i)})=Y^{2}(Z^{\prime}_{\mathcal{N}_{G^{\prime}}(i)})\qquad$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
@@
A similar approach to defining non-sharp null of spillovers in the literature is through the use of exposure functions (Aronow and Samii (2017); Basse et al. (2019b); Puelz et al. (2022)). The exposure functions are functions of the treatment assignment vector, $Z$, such that restriction sets can be created on $Z$ by partitioning $Z$ on the range of the exposure function. In the above Example 1, we can define an exposure function as below.
@@
We obtain the setup described in the literature, given a fixed $G$ using the above-described exposure function as a restriction set. The setup assumes a spillover mechanism via an exposure function. Since we consider the graph underlying the population to be random, we define restrictions on the joint treatment assignment in Hypothesis 3. A spillover mechanism, if existing, will be captured in the potential outcome function, and hence, we do not separately define a joint exposure mapping function by generalizing the exposure mappings. Careful identification of a spillover mechanism will be required for the inversion of the hypothesis test to obtain the confidence interval. We describe the currently proposed method in the literature of a conditional-randomization test procedure to obtain valid p-values (Puelz et al. (2022)).
Procedure: Consider $C$ to be a subset of units and treatment assignment $Z$ such that a test statistic $T(Z|C)$ is imputable under the null.
@@
The above-stated procedure holds finite-sample validity. We note that the conditioning mechanism $C$ is drawn from $P(C|Z_{obs})$, which should be determined by the analyst. One must ensure the imputability of the test statistic $T(Z|C)$ when constructing a mechanism for $C$. We require the test statistic to be imputable to obtain its sampling distribution under the null, even though we observe only one treatment assignment vector. Below, we define the imputability of a test statistic for the joint treatment assignment.
@@
3   Quasi-randomization test based on network distribution
We present the quasi-randomization test for the sharp null of no spillover effects. Our main idea is to generate the null sampling distribution of a test statistic by randomizing over the network-generating process. Intuitively, under the null of no spillover effect, the underlying network structure does not contribute to the total treatment effect. This can also be seen in Hypothesis 4 where the network variable $G$ is unrestricted in the, otherwise, non-sharp null. This leads to the imputability of a test statistic of choice by fixing the direct treatment assignment $Z$ and randomizing over the distribution of network $G$. We formally define the conditional test statistic, $T_{c}$, below and show its imputability under the null Hypothesis 4.
@@ 3   Quasi-randomization test based on network distribution
$Y(Z,G)=Y(Z^{\prime},G^{\prime})\qquad Z=Z^{\prime},$
@@
$=T(Z_{obs},G,Y(Z_{obs},G^{\prime}))\qquad P(G^{\prime})>0\text{ under }\mathcal{H}_{0}.$
@@
We state our proposed quasi-randomization testing procedure below. The procedure assumes the knowledge of the network-generating process $P(G)$. An elaborate consideration of the distribution of the network will be presented in Section 4.
@@
Consider the null hypothesis, $\mathcal{H}_{0}$, in Hypothesis 4. Let $Z_{obs}\sim P(Z)$ be the treatment assignment of a randomized experiment. We assume that $G_{obs}\sim P(G)$ where $P(G)$ is known. Let $T(G|Z)$ be an imputable test statistic under the null $\mathcal{H}_{0}$. Then, Procedure 1 described above is conditionally valid at level $\alpha\in(0,1)$. That is,
$\mathbb{E}\bigl{[}\mathcal{I}(\text{p-val}(Z_{obs},G_{obs},Y_{obs})\leq\alpha)|\mathcal{H}_{0}\bigl{]}\leq\alpha\qquad\forall\text{ }\alpha\in(0,1),$
Here, the expectation is with respect to the distribution of $P(G_{obs}|Z_{obs})$.
@@
Note that conditional validity implies unconditional validity of the procedure too. A crucial component of the proof is the identification of the conditional distribution $P(G|Z)$ for the validity of Procedure 1. This is contingent upon the experimental design, and different experimental designs will lead to different conditional randomization distributions. We compare this with the state-of-the-art methodology presented in Section 2 where the null randomization distribution comes from $P(Z|C)$. Here, $C$ is a conditioning mechanism consisting of a subset of chosen units. Thus, different experimental designs lead to different $P(Z|C)$. A common problem faced in the literature is in the implementation of complex experimental designs like cluster randomized experiments. In a cluster-randomized experiment, the population is classified into a set of clusters where treatments are assigned at the cluster level. If a cluster is treated, then all units in that cluster receive the treatment or vice-versa. Fixing the treatment assignment of units in the conditioning mechanism might render the null randomization distribution $P(Z|C)$ degenerate, given the restricted span of cluster treatment assignments space. For example, assume there is a focal unit in every cluster. Then, conditioning on the observed treatment of focal units, we obtain $Z|C=\{Z_{obs}\}$ since all the unit treatment assignments can be deterministically imputed by the treatment assignment of the focal unit in that cluster. This reduction in the support of conditional distribution leads to either degeneracy or enormous loss of power. Our methodology captures the null distribution by randomizing over $G|Z$, where the null distribution admits large support. We discuss this in Sections 3.1.1 and 3.1.2, where we identify the null randomization distribution under different experimental designs.
@@ 3.1.1   Completely randomized experiment
$P(Z=z)=\begin{cases}\frac{1}{\binom{N}{N_{t}}}&\sum_{i=1}^{N}z_{i}=N_{t},\\ 0&\text{otherwise}.\end{cases}$ (1)
We remark that $G\!\perp\!\!\!\perp Z$ in Equation 1. Hence, for completely randomized experiments,
@@ 3.1.2   Cluster randomized experiment
4   Quasi-permutation test via distributional properties
Identification of null randomization distribution $P(G|Z)$ involves knowledge of the distribution of the network-generating process, which is not identified by the design of the experiment. To this end, we explore random graph models to model $P(G)$. As discussed in the previous section, this identification of null randomization distribution becomes challenging when the complexity of the experimental design increases. This challenge is overcome in our proposed methodology, which is a conditional permutation test. The method does not involve estimating the null distribution. We discuss this in Section 4.1.2.
@@ 4.1   Random graph models
In this section, we describe various random graph models based on which $P(G)$ can be modeled. The seminal paper by Erdős and Rényi Erdős et al. (1960) introduced the concept of random graph for the first time. The Erdős-Rényi random graph model has since been studied widely, owing to exact computations of desirable graph properties. We now present the model formally in Section 4.1.1. However, the Erdős-Rényi random graph model does not represent most real-world networks, such as social networks, supply-chain networks, and scientific collaboration networks, among many others. We look at a generalization of Erdős-Rényi random graph that is better adept at modeling such real-world networks, proposed by Newman et al. (2002), in Section 4.1.2.
4.1.1   Erdős-Rényi random graph model
We formally describe the Erdős-Rényi random graph model (Erdős et al. (1960)). The below-described model is the probabilist variant of the (Erdős-Rényi) random graph model. We consider the number of nodes, $N$, and edges, $|E|$, to be fixed, and any two nodes will be linked independently with an edge with a fixed probability $p$. That is,
In order to model $P(G)$ using Equation (3), we have to estimate the parameter $p$. We obtain the MLE of the parameter $p$ and estimate it as $\frac{|E|}{\binom{N}{2}}$. We conduct an extensive numerical simulation study and verify the finite sample validity of the proposed randomization test. We also obtain the power of the test under different scenarios, which are presented in Appendix C. An Erdős-Rényi random graph model is the simplest network model one can consider, and it may not be representative of the kind of networks we observe in the real world. For example, in social networks, one observes a high degree of clustering in the graph (Watts and Strogatz (1998)). One drawback of the model is the independence and identical probability of formation of two edges in the graph. One way to characterize this is to evaluate the degree distribution of the random graph. It has been shown that the degree distribution of a large Erdős-Rényi random graph is a Poisson distribution. Most real-life networks violate this property and have highly skewed degree distribution. In the next section, we generalize this and characterize random graphs with arbitrary degree distribution.
@@ 4.1.2   Random graphs with arbitrary degree distribution
A key feature of random graphs with arbitrary degree distribution is their ability to capture a large class of degree distribution. This generalizes the Erdős-Rényi random graph model and, at the same time, ensures tractability of important structural graph properties (Newman et al. (2002)). In order to deploy Procedure 1, we have to identify $P(G)$. This involves estimating $(p_{1},p_{2},...,p_{N})$ with a single realization of the graph, and we cannot directly implement Procedure 1 without estimating $P(G)$. In the case of random graphs with arbitrary degree distribution, it is proven that we can possibly estimate the parameters of the random graph with just one realization but with some uncertainty (Chatterjee et al. (2011)). Owing to estimation issues of the null distribution, we propose a new method that does not involve such estimations and reduces to a simple permutation test. In our proposed methodology, we use sufficient statistics of the random graph distribution, degree sequence in this case, and develop a method conditioning on the sufficient statistic. We use the symmetric nature of the $P(G)$ via the equivalence relation and build a conditional permutation test, as described below.
@@ 4.2   Permutation test based on equiprobability events
Here, we present our main result, which proposes the procedure to obtain a conditional randomization test without estimating the distribution. Our main idea is to identify a sufficient statistic for the random graph model, which is degree sequence in the case of random graphs with arbitrary degree distribution. We then sample the null distribution over all the graphs with the same degree sequence as the observed graph. Using the equiprobability of the event set of all the graphs with the same degree sequence, we obtain a permutation test. We now formally present the definitions below and proceed to state and prove the finite sample validity of the proposed method.
@@
Consider $D=(d_{1},d_{2},...,d_{N})$ such that $d_{1}\leq d_{2}\leq...\leq d_{N}$. We call $D$ to be the degree sequence of $G=(V,E)$ if there exists a one-to-one mapping from $\{deg(i):i\in V\}$ to $D$. We denote $D$ as $deg(G)$.
Not all $D\in\mathbb{W}^{N}$ can be a valid degree sequence. For example, consider $D=\{1,2\}$ as a candidate for degree sequence for a graph of size $2$. We remark that the maximum degree of a node in the graph is at most $(N-1)$. Since no vertex in a graph of size two can have a degree of more than 1, $D$ is not a valid degree sequence. Note that graphs with the same degree sequence are not necessarily isomorphic graphs. This implies that the vertices of the graph cannot be permuted to obtain the other graph with edges preserved (see Figure 1).
@@
The converse holds true. Consider a set of distinct graphs with the number of units fixed to $N$ and a degree sequence of the graphs fixed to a valid degree sequence $D$. Then, this set of all graphs with degree sequence $D$ will form an equivalence class, and we can partition the space of all random graphs with this equivalence relation. We define this below.
@@
It is a quick check that $d$ defines an equivalence relation. Hence, we can partition the support of $P(G)$, where each part in the partition represents an equivalence class of $D$.
@@
Consider the null hypothesis of no spillover effects in Hypothesis 4, denoted by $\mathcal{H}_{0}$. Let $Z_{obs}\sim P(Z)$ be the treatment assignment of a cluster randomized experiment. We assume that $G_{obs}\sim P(G)$ where $G$ is a random graph with arbitrary degree distribution. Let $T(G|Z)$ be an imputable test statistic under $\mathcal{H}_{0}$. Then, Procedure 2 described above is conditionally valid at level $\alpha\in(0,1)$. That is,
$\mathbb{E}\bigl{[}\mathcal{I}(p(Z_{obs},G_{obs},Y_{obs})\leq\alpha)|\mathcal{H}_{0}\bigl{]}\leq\alpha\qquad\forall\text{ }\alpha\in(0,1),$
Here, the expectation is taken with respect to the distribution of $P(G_{obs}|Z_{obs})$.
@@
As a corollary, one can take all the clusters to be of size one, implying every unit is a cluster. This retrieves the setup for a completely randomized experiment, and thus, Theorem 2 also holds true for a completely randomized experiment.
We emphasize that the generation of graphs with a given degree sequence admits a polynomial time algorithm. Given fixed parameters like the maximum degree of the graph, one can use state-of-the-art algorithms to efficiently sample graphs with a given degree sequence uniformly from the space of all the graphs with the same degree sequence (Arman et al. (2021)). It has been shown that the size of this equivalence class can be exponentially large (Barvinok and Hartigan (2013)). It should be noted that one should use a uniform sampler in order to avoid introducing any bias in the null distribution attributed to sampling. We can also overcome this by taking a large number of samples to approximate the null distribution effectively.

### v2

@@ 2.3 Null hypothesis of no spillover and conditional randomization test
$\forall i\in[N]\quad Y_{i}(Z,G)=Y_{i}(Z^{\prime},G^{\prime})$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
@@
$\forall i\in[N]\quad Y^{2}(Z_{\mathcal{N}_{G}(i)})=Y^{2}(Z^{\prime}_{\mathcal{N}_{G^{\prime}}(i)})$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
@@
This approach of defining non-sharp null of spillovers in the literature is through the use of exposure functions (Aronow and Samii (2017); Basse et al. (2019b); Puelz et al. (2022)). The exposure functions are functions of the treatment assignment vector, $Z$, such that restriction sets can be created on $Z$ by partitioning $Z$ on the range of the exposure function. We can define an exposure function in the above Example 1:
@@
We obtain the setup described in the literature, given a fixed $G$ using the above-described exposure function as a restriction set. The setup assumes a spillover mechanism via an exposure function. Since we consider the graph underlying the population to be random, we define restrictions on the joint treatment assignment in Hypothesis 3. A spillover mechanism, if existing, will be captured in the potential outcome function. We describe the currently proposed method in the literature of a conditional-randomization test procedure to obtain valid p-values (Puelz et al. (2022)).
Procedure 0: Consider $C$ as a subset of units and treatment assignment $Z$ such that a test statistic $T(Z|C)$ is imputable under the null.
@@
The above-stated procedure holds finite-sample validity. The conditioning mechanism $C$ is drawn from $P(C|Z_{obs})$, which the analyst should determine. One must ensure the imputability of the test statistic $T(Z|C)$ when constructing a mechanism for $C$. We require the test statistic to be imputable to obtain its sampling distribution under the null, even though we observe only one treatment assignment vector. Below, we define the imputability of a test statistic for the joint treatment assignment.
@@
3 Quasi-randomization test based on network
We present the quasi-randomization test for the sharp null of no spillover effects. Our main idea is to generate the null sampling distribution of a test statistic by randomizing over the network-generating process. Intuitively, the underlying network structure does not contribute to the total treatment effect under the null of no spillover effect. This can also be seen in Hypothesis 4 where the network variable $G$ is unrestricted in the, otherwise, non-sharp null. This leads to the imputability of a test statistic of choice by fixing the direct treatment assignment $Z$ and randomizing over the network distribution $G$. We formally define the conditional test statistic, $T_{c}$, below and show its imputability under the null Hypothesis 4.
@@ 3 Quasi-randomization test based on network
$Y(Z,G)=Y(Z^{\prime},G^{\prime})\qquad Z=Z^{\prime}.$
@@
$=T(Z_{obs},G,Y(Z_{obs},G^{\prime}))\qquad P(G^{\prime}|Z_{obs})>0\text{ under }\mathcal{H}_{0}.$
@@
We state our proposed quasi-randomization testing procedure below. The procedure assumes the knowledge of the network-generating process $P(G)$. An elaborate consideration of the network distribution will be presented in Section 4.
@@
Consider the null hypothesis, $\mathcal{H}_{0}$, in Hypothesis 4. Let $(Z_{obs},G_{obs})\sim P(Z,G)$ be the bivariate treatment assignment of a randomized experiment. We assume that $P(Z,G)$ is known. Consider a test statistic $T(Z,G,Y(Z,G))$. Then, Procedure 1 described above is conditionally valid at level $\alpha\in(0,1)$. That is,
$\mathbb{E}\bigl{[}\mathcal{I}(\text{p-val}(Z_{obs},G_{obs},Y_{obs})\leq\alpha)|\mathcal{H}_{0}\bigl{]}\leq\alpha\qquad\forall\text{ }\alpha\in(0,1).$
Here, the expectation is for the distribution of $P(G_{obs}|Z_{obs})$.
@@
Note that conditional validity implies unconditional validity of the procedure, too. A crucial component of the proof is identifying the conditional distribution $P(G|Z)$ for the validity of Procedure 1. This is contingent upon the experimental design, and different experimental designs will lead to different conditional randomization distributions. We compare this with the state-of-the-art methodology presented in Section 2 where the null randomization distribution comes from $P(Z|C)$. Here, $C$ is a conditioning mechanism consisting of a subset of chosen units. Thus, different experimental designs lead to different $P(Z|C)$. A common problem faced in the literature is implementing the methods in cluster randomized experiments. In a cluster-randomized experiment, the population is classified into clusters where treatments are assigned at the cluster level. If a cluster is treated, all units in that cluster receive the treatment or vice-versa. Fixing the treatment assignment of units in the conditioning mechanism might render the null randomization distribution $P(Z|C)$ degenerate, given the restricted span of cluster treatment assignments space. For example, assume there is a focal unit in every cluster. Then, conditioning on the observed treatment of focal units, we obtain $Z|C=\{Z_{obs}\}$ since all the unit treatment assignments can be deterministically imputed by the treatment assignment of the focal unit in that cluster. This reduction in the support of conditional distribution leads to either degeneracy or enormous loss of power. Our methodology captures the null distribution by randomizing over $G|Z$, where the null distribution admits ample support. We discuss this in Sections 3.1.1 and 3.1.2, where we identify the null randomization distribution under different experimental designs.
@@ 3.1.1 Completely randomized experiment
$P(Z=z\;|\;G)=\begin{cases}\frac{1}{\binom{N}{N_{t}}}&\sum_{i=1}^{N}z_{i}=N_{t},\\ 0&\text{otherwise}.\end{cases}$ (1)
We remark that $G\!\perp\!\!\!\perp Z$ by design. Hence, for completely randomized experiments,
@@ 3.1.2 Cluster randomized experiment
4 Permutation test via distributional properties
Identification of null randomization distribution $P(G|Z)$ involves knowledge of the distribution of the network-generating process, which is not identified by the design of the experiment. To this end, we explore random graph models to model $P(G)$. As discussed in the previous section, identifying null randomization distribution becomes challenging in cluster-randomized experiments. This challenge is overcome in our proposed methodology, which is a conditional permutation test. We discuss this in Section 4.1.1.
@@ 4.1 Random graph models
This section describes various random graph models based on which $P(G)$ can be modeled. The Erdős-Rényi random graph model has been studied widely, owing to exact computations of desirable graph properties. We present the model formally in Appendix C.1. However, the Erdős-Rényi random graph model does not represent most real-world networks, such as social networks, supply-chain networks, and scientific collaboration networks, among many others. We look at a generalization of Erdős-Rényi random graph better at modeling such real-world networks, proposed by Newman et al. (2002), in Section 4.1.1. For this article, we restrict the models to be covariate-free.
@@ 4.1.1 Random graphs with arbitrary degree distribution
A key feature of random graphs with arbitrary degree distribution is their ability to capture a large class of degree distribution. This generalizes the Erdős-Rényi random graph model and, at the same time, ensures tractability of important structural graph properties (Newman et al. (2002)). To deploy Procedure 1, we must identify $P(G)$. This involves estimating $(p_{1},p_{2},...,p_{N})$ with a single realization of the graph, and we cannot directly implement Procedure 1 without estimating $P(G)$. In the case of random graphs with arbitrary degree distribution, it is proven that we can estimate the parameters of the random graph with just one realization but with some uncertainty (Chatterjee et al. (2011)). However, there might be other relational dependencies characterizing the formation of the network that are not captured by random graphs with arbitrary degree distribution. For example, the phenomena of transitivity are not captured in this model, where two nodes have higher chances of being connected if they have a common connecting node.
4.1.2 Exponential random graph model
The state-of-the-art method for characterizing random graph formation of social networks is through the exponential random graph model or ERGM (Snijders et al. (2006)). It is a generalization of the random graphs with arbitrary degree distribution, where the characterization of the distribution of the random graph is through a set of given sufficient statistics. This could also contain information about covariates to capture homophily in the network formation process. In this article, we consider the random graph formation free of covariates.
Definition 4.
Consider a random graph $G\in\{0,1\}^{N\times N}$ such that $P(G=g|\eta)=\frac{1}{\kappa}\cdot{e^{\eta^{T}\cdot s(g)}}$. Here, $s(g)$ represents a vector of sufficient statistics, and $\kappa$ is the normalizing constant. We call $G$ an exponential random graph.
Common sufficient statistics used in the literature for the characterization of social networks include the number of edges, the number of triangles, and the number of stars of different sizes, say stars of size two or stars of size three (Robins et al. (2007)). It is important to carefully choose the model as an ill-posed ERGM can lead to near degeneracy (Chatterjee and Diaconis (2013); Schweinberger et al. (2020)). As with random graphs with arbitrary degree distribution, estimating ERGM is fraught with challenges given just a single realization.
Owing to estimation issues of the null distribution, we propose a new method that does not involve such estimations and reduces to a simple permutation test. In our proposed methodology, we first consider the sufficient statistics of the random graph distribution as degree sequence and develop a method conditioning on the sufficient statistic. We then generalize the method to a set of arbitrary sufficient statistics. We use the symmetric nature of the $P(G)$ via an equivalence relation and build a conditional permutation test, as described below.
@@ 4.2 Permutation test based on equiprobability events
4.2.1 Test for a random graph with arbitrary degree distribution
Here, we present our main result, which proposes obtaining a conditional randomization test without estimating the distribution. Our main idea is to identify a sufficient statistic for the random graph model, which is degree sequence in the case of random graphs with arbitrary degree distribution, and condition on this sufficient statistic. We sample the null distribution over all the graphs with the same degree sequence as the observed graph. We obtain a permutation test using the equiprobability of the event set of all the graphs with the same degree sequence. We now formally present the definitions below and proceed to state and prove the finite sample validity of the proposed method.
@@ 4.2.1 Test for a random graph with arbitrary degree distribution
Consider $D=(d_{1},d_{2},...,d_{N})$ such that $d_{1}\leq d_{2}\leq...\leq d_{N}$. We call $D$ to be the degree sequence of $G=(V=\{v_{1},v_{2},...,v_{N}\},E)$ if there exists a one-to-one mapping $\delta_{G}:V\rightarrow D$ such that $\delta(v_{i})=deg(v_{j})$ for some $v_{j}\in V$ for all $v_{i}\in V$. We denote $D$ as $deg(G)$.
Not all $D\in\mathbb{W}^{N}$ can be a valid degree sequence. For example, consider $D=\{1,2\}$ as a candidate for degree sequence for a graph of size $2$. We remark that the maximum degree of a node in the graph is at most $(N-1)$. Since no vertex in a graph of size two can have a degree of more than 1, $D$ is not a valid degree sequence. Note that graphs with the same degree sequence are not necessarily isomorphic graphs. This implies that the graph’s vertices cannot be permuted to obtain the other graph with edges preserved (see Figure 1).
@@
The converse holds. Consider a set of distinct graphs with the number of units fixed to $N$ and a degree sequence of the graphs fixed to a valid degree sequence $D$. Then, this set of all graphs with degree sequence $D$ will form an equivalence class, and we can partition the space of all random graphs with this equivalence relation. We define this below.
@@
It is a quick check that $d$ defines an equivalence relation. Hence, we can partition the support of $P(G)$, where each part represents an equivalence class of $D$.
@@
Consider the null hypothesis of no spillover effects in Hypothesis 4, denoted by $\mathcal{H}_{0}$. Let $(Z_{obs},G_{obs})\sim P(Z,G)$ be the bivariate treatment assignment of a completely randomized experiment. We assume that $G_{obs}\sim P_{G}$ where $G$ is a random graph with arbitrary degree distribution. Consider a test statistic $T(Z,G,Y(Z,G))$. Then, Procedure 2 described above is conditionally valid at level $\alpha\in(0,1)$. That is,
$\mathbb{E}\bigl{[}\mathcal{I}(p(Z_{obs},G_{obs},Y_{obs})\leq\alpha)|\mathcal{H}_{0}\bigl{]}\leq\alpha\qquad\forall\text{ }\alpha\in(0,1).$
Here, the expectation is taken for the distribution of $P(G_{obs}|Z_{obs})$.
@@
As a corollary, we can substitute any sub-class of the degree sequence equivalence class in Procedure 2, and the finite-sample validity of the Procedure is retained. For example, we can take the equivalence class to be all the graphs isomorphic to the observed graph in the second step of Procedure 2. We present results for both these Procedures in Section 5.
We emphasize that generating graphs with a given degree sequence admits a polynomial time algorithm. Given fixed parameters like the maximum degree of the graph, one can use state-of-the-art algorithms to efficiently sample graphs with a given degree sequence uniformly from the space of all the graphs with the same degree sequence (Arman et al. (2021)). It has been shown that the size of this equivalence class can be exponentially large (Barvinok and Hartigan (2013)). It should be noted that one should use a uniform sampler to avoid introducing any bias in the null distribution attributed to sampling. We can also overcome this by effectively taking many samples to approximate the null distribution.
4.2.2 Test for exponential random graph model
We generalize the above method to a class of exponential random graph models where the sufficient statistics of the model are permutation invariant. We achieve this by conditioning on the graph isomorphism class and generating the null cases from this equivalence class. We formally present the results below.
A graph is considered isomorphic to another graph if a one-to-one correspondence between the vertices of the two graphs preserves the edges between the graphs. Below, we present an equivalence condition for the same.
Definition 7.
Two graphs $G$ and $G^{\prime}$ are isomorphic if and only if a permutation matrix $P$ exists such that $G=P^{T}G^{\prime}P$. We denote $G$ and $G^{\prime}$ as isomorphic graphs using $G\cong G^{\prime}$.
We note that graph isomorphism defines an equivalence relation. The permutation matrix comprises row binary vectors indexed by the one-to-one correspondence between the vertices of the graphs. We now define a class of exponential random graph models based on the permutation invariance property of its sufficient statistic.
Definition 8.
An exponential random graph model $G\in\{0,1\}^{N\times N}$ is called permutation-invariant if $s(\Pi(g))=s(g)$. Here, $\Pi$ represents the permutation function of the graph’s vertices, and $s(.)$ represents the vector of sufficient statistics of the exponential random graph model.
Many sufficient statistics used in real-life social network modeling follow this property. As discussed, some graph statistics include the number of triangles, degree sequence, and number of stars of size two, and they are permutation invariant. We model the data-generating process as an exponential random graph with permutation invariance property and present a permutation testing procedure.
Procedure 3: Let $(Z_{obs},G_{obs},Y_{obs})$ be the joint treatment assignment vector and its corresponding outcome. Consider a test statistic $T_{c}(G|Z_{obs}):=T(Z_{obs},G,Y_{obs})$.
  1. 1.
  2. 2.
    Consider the sets $Z^{t}:=\{i\in\mathbb{P}:\>Z_{i}=t\}\;\forall\;t\in\{0,1\}$.
  3. 3.
    Define the equivalence class
    $[G_{obs}|Z_{obs}]=\{G\in\mathcal{G}\;|\>G[{Z_{obs}}^{0}]\cong G_{obs}[{Z_{obs}}^{0}],\>G[{Z_{obs}}^{1}]\cong G_{obs}[{Z_{obs}}^{1}]\}.$
  4. 4.
    Compute p-value as
    $p(Z_{obs},G_{obs},Y_{obs}):=\frac{1}{|[G_{obs}|Z_{obs}]|}\cdot\sum_{G\in[G_{obs}|Z_{obs}]}\mathcal{I}(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})).$
Theorem 3.
Consider the null hypothesis of no spillover effects in Hypothesis 4, denoted by $\mathcal{H}_{0}$. Let $(Z_{obs},G_{obs})\sim P(Z,G)$ be the bivariate treatment assignment of a cluster randomized experiment. We assume that $G_{obs}\sim P_{G}$ where $G$ is a permutation-invariant exponential random graph model. Consider a test statistic $T(Z,G,Y(Z,G))$. Then, Procedure 3 described above is conditionally valid at level $\alpha\in(0,1)$. That is,
$\mathbb{E}\bigl{[}\mathcal{I}(p(Z_{obs},G_{obs},Y_{obs})\leq\alpha)|\mathcal{H}_{0}\bigl{]}\leq\alpha\qquad\forall\text{ }\alpha\in(0,1).$
Here, the expectation is taken for the distribution of $P(G_{obs}|Z_{obs})$.
Proof.
Refer to Appendix B.2. ∎